OpenStudy (yanasidlinskiy):

Answers that need to be checked! 1.) Determine whether each ordered pair is a solution of the system of linear equations. 3y=9x+12 3x-y=-4 a. (-2,-2) b. (0,4) MY ANSWER: I PUT YES

OpenStudy (yanasidlinskiy):

2.) Solve the system of equations by graphing. x-y=4 y=-1 MY ANSWER: (1,-3) (0,-4)

OpenStudy (yanasidlinskiy):

Lines intersect @ (1, -3)

Nnesha (nnesha):

yep that's right but why did you put 0 , -4 for 2nd question ?

OpenStudy (yanasidlinskiy):

3.) Solve the system of equations by graphing. y=-4x+2 3y+12x=6 Lines intersect @ (1, -2)

OpenStudy (yanasidlinskiy):

I have to find 2 points to plot it on the graph.

OpenStudy (yanasidlinskiy):

Solve the system of equations by substitution. x=7y+6 -2x+14y=-12 I got infinitely many solutions.

Nnesha (nnesha):

last one when you solve 2nd equation for y you will get 1st equation so i would write infinitely many solution

OpenStudy (lurker):

1)b 2) wrong 3)right 4)right

Nnesha (nnesha):

i mean 3rd question

OpenStudy (yanasidlinskiy):

I'm confused....

OpenStudy (yanasidlinskiy):

What do you mean?

OpenStudy (yanasidlinskiy):

Are you saying I solved for number 3 wrong then?

Nnesha (nnesha):

what's the statement for 3rd question ? they are asking 4 solution ?

OpenStudy (yanasidlinskiy):

For number 3, they are asking my to solve by graphing. Y HAS ALREADY been solved for.. Now I need to solve for the other equation. When solving for the other equation and plotting it on the graph, lines intersect at points (1, -2). Is the intersection point correct?

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @YanaSidlinskiy 3.) Solve the system of equations by graphing. y=-4x+2 3y+12x=6 Lines intersect @ (1, -2) \(\color{blue}{\text{End of Quote}}\) for this one you will get same line both equations are same so thats mean they are intersecting at every point

Nnesha (nnesha):

solve 2nd equation for y \[\huge\rm 3y+12x=6\] subtract 12x both sides and then divide by 3

Nnesha (nnesha):

its called dependent system when 2 equations represent the same line (same slope ,same y-intercept )

OpenStudy (yanasidlinskiy):

This topic is on solving by the graphing method....

Nnesha (nnesha):

okay so when you graph both equation you will get same line there isn't ONE point where they intersect so answer would be infinitely many solution

Nnesha (nnesha):

1) right 2nd) you just need a point where both lines intersect ,not the y-intercept 4th) right

Nnesha (nnesha):

let me know if u r still confused :=)

OpenStudy (yanasidlinskiy):

Thanks! Glad you could help out a little!